# NeuroTree: Hierarchical Functional Brain Pathway Decoding for Mental Health Disorders

### v2

@@ 6 Experiments
In this section, we first evaluate the proposed NeuroTree performance in disease classification (section 6.3) and its application to predicting chronological age (section 9). In section 7.2, we analyze the convergence of the spectral norm for $k$-hop connectivity. Finally, we examine how hierarchical brain trees interpret FC patterns between brain regions and their corresponding seven subnetworks in section 8.
Table 2: Summary statistics of demographics in Cannabis and COBRE
Dataset Sample size Age (Mean ± SD)
Cannabis HC (n=128) 30.06 $\pm$ 10.47
Cannabis (n=195) 27.23 $\pm$ 7.73
COBRE HC (n=72) 38.31 $\pm$ 12.21
Schizophrenia (n=70) 36.04 $\pm$ 13.02
@@ 6.1 Datasets
We validated two publicly available fMRI datasets —one focusing on cannabis use disorder and the other on schizophrenia. ① Cannabis [12]: The cannabis dataset comprises fMRI data from two distinct sources. The data were preprocessed from 3-Tesla fMRI acquisitions, and the mean time series for each subject was computed across 90 ROIs using the Stanford atlas parcellation [46]. ② COBRE [47]: The Center for Biomedical Research Excellence (COBRE) dataset includes resting-state fMRI data collected from healthy controls and individuals diagnosed with schizophrenia. All MRI data were parcellated into 118 ROI regions using the Harvard-Oxford atlas. We summarized the statistics and demographics of both datasets in Table 2.
Comparison Models: Classical methods for constructing brain networks typically rely on Pearson correlation and $k$-nearest neighbor approaches combined with GCNs. The baseline methods evaluated include) the $Pearson$ GCN: This model constructs an fMRI graph using Pearson correlation coefficients to represent relationships between ROIs. 2) $k$-NN GCN: This model builds brain networks by connecting k-nearest neighbors based on similar input features among brain regions. 3) GAT [33]: This model extends the GCN framework by incorporating an attention mechanism that enhances feature aggregation and learning from neighboring nodes. Addictionally, we also evaluating the performance of NeuroTree against state-of-the-art (SOTA) GCN architectures including 4) BrainGNN [41]: BrainGNN is an interpretable GNN framework designed for analyzing fMRI data. This framework effectively identifies and characterizes significant brain regions and community patterns associated with specific neurological conditions. 5) BrainUSL [42]: BrainUSL learns the graph structure directly from BOLD signals while incorporating sparsity constraints and contrastive learning to capture meaningful connections between brain regions. Finally, we compare NeuroTree with other methods that incorporate higher-order node information, specifically MixHop [44] and graph path-based models such as GPC-GCN [45] and the tree path-based model PathNN [31].
6.2 Experiments setting
In our experimental setup, we employed the following training parameters: scale parameter ($\rho$) of 0.1, 2-time segments ($T$), and a $k$-hop connectivity balance ($\lambda$) of 0.5 between dynamic FC matrices, and multi-head with four heads. We use Area Under the Curve (AUC), Accuracy (Acc.), Precision (Prec.), and Recall (Rec.) as metrics for graph classification, as well as the Mean Squared Error (MSE) metric for the regression task for age prediction. The comparative analysis was conducted using 5-fold cross-validation on two distinct datasets, benchmarking NeuroTree against baseline models and SOTA models.
@@ 6.2 Experiments setting
Figure 3: Convergence analysis of $\Phi_{k}(t)$ over $k$-hop The spectral norm of $\Phi_{k}(t)$ exhibits differential convergence rates across varying $k$-orders among different subtyping diseases. The cannabis dataset demonstrates a steeper convergence gradient than COBRE as the regularization parameter $\lambda$ increases.
6.3 Compare with Brain GCN Models
@@
We quantitatively assessed brain network construction using various metrics, employing Pearson correlation and KNN as connectivity measures. As shown in Table 1, the GAT model effectively integrates attention mechanisms into fMRI node features, achieving mean AUC scores of 0.72 and 0.67. In contrast, the conventional GCN is limited to learning solely from topological structures. Furthermore, NeuroTree enhances ACC performance in fMRI-based models, improving BrainGNN (ACC $\uparrow$ 27.11%), BrainUSL (ACC $\uparrow$ 15.38%), and BrainGSL (ACC $\uparrow$ 15.38%) on the cannabis dataset. On the COBRE dataset, BrainGNN, BrainUSL, and BrainGSL further improve ACC by 38.00%, 27.77%, and 35.29%, respectively.
6.4 Compare with Path-Based Models
@@
As shown in Table 1, NeuroTree outperforms existing high-connectivity view field-aware graph GCN models, such as the high-order concatenation layer-based MixHop, achieving over 5% average performance improvement in AUC across all metrics on the cannabis dataset. Furthermore, compared to MixHop on the COBRE dataset, NeuroTree not only achieved a 2.8% increase in AUC, but also demonstrated a mean performance improvement of 13.1%, 11.3%, and 11.4% in ACC, Precision, and Recall. Notably, although GPC-GCN can aggregate multi-level connectivity information by considering the concept of paths, it does not significantly improve classification performance on critical connectivity paths. Similarly, the PathNN framework aggregates path information starting from nodes to generate node representations but shows only limited improvements. In contrast, our NeuroTree achieves a 10% improvement in AUC on the cannabis dataset, while PathNN underperforms compared to MixHop on the COBRE dataset. The experimental results indicate that NeuroTree is superior in capturing temporal functional connectivity patterns and predicting outcomes in addiction and psychiatric disorders.
@@
Figure 4: The brain tree visualization illustrates psychiatric disorders structured into three hierarchical trunk levels. Panels (a-1) and (b-1) highlight the most significant nodes along the tree path. The $l$-1 pathways represent regional connectivity, corresponding to the level three brain maps on the right. Panels (a-2) and (a-3) depict the number of connections using color gradients across hierarchical levels.
7 Ablation Study
7.1 The Impact of Objective Function and Demographics
@@
In this section, we evaluate the efficacy of CMF loss integration within our proposed NeuroTree. The quantitative analysis in Table 1 demonstrates that without $\mathcal{L}_{CMF}$, the model only achieves an AUC of 0.74 on the cannabis dataset and 0.63 on the COBRE dataset. Notably, even without the CMF loss component, our approach still exhibited performance comparable to MixHop across other metrics, particularly in characterizing mental disorder feature classification. With CMF loss integration, NeuroTree achieves superior results with AUC values of 0.77 and 0.71, respectively. This enhancement stems from combining the loss function with an attention mechanism optimization framework that simultaneously suppresses connectivity between dissimilar regions while amplifying patterns among similar neurological areas. Furthermore, our ablation experiments demonstrate that age serves as a critical modulator in AGE-GCN by adapting parameter $\theta$, substantially influencing the classification of changes in brain connectivity.
@@ 7.2 Convergence Analysis
To further examine the convergence speed of $k$-hop FC in various brain disorders, we analyze the spectral norm under the $k=1$ setting in the convergence of $\Phi_{k}(t)$. Although the COBRE dataset initially exhibits slower convergence, particularly with the boundary defined around order 15, the overall convergence in the cannabis dataset is faster than in COBRE across different $k$-hop values. Furthermore, when analyzing the convergence of $\Phi_{k}(t)$ under varying $\lambda$ values, as shown in Fig. 3 (middle) and (right), the $\lambda$ values for cannabis are more concentrated and stable. In contrast, COBRE requires larger $\lambda$ values to achieve convergence. The empirical results effectively capture the relationship between dynamic and static FC of $\hat{A}_{k}(t)$. This indicates that addiction may lead to stronger brain connectivity signals compared to psychiatric disorders, as reflected in the affected brain regions.
Figure 5: Average FC strength weights versus $\alpha$ in high-order tree paths, comparing (a) HC versus cannabis users and (b) HC versus schizophrenia patients. The results indicate reduced connectivity patterns with increasing $\alpha$ in both clinical groups compared to the control group.
Figure 6: Heatmaps illustrate connectivity alterations across Yeo’s seven networks during age progression in (a) addiction and (b) schizophrenia. Below, each age stage maps the seven subnetworks in both hemispheres, shown in medial and lateral views.
@@ 8 Brain Tree Analysis
This section investigates brain disorders by analyzing hierarchical brain networks using decoded fMRI-derived connectomes. NeuroTree deconstructs the brain network into a tree-structured component graph consisting of a main trunk and radial branches. We identify key regions associated with brain disorders within each trunk by analyzing optimal weighted high-order tree paths. Our analysis highlights the most relevant functional network regions determining critical pathways within the brain connectome tree.
8.1 Exploring Hierarchical Regional Patterns in Brain Disorders
@@
To better understand how embedded brain features vary across different hierarchical levels of brain subnetworks, NeuroTree predicts each ROI score and performs reranking, as illustrated in Fig. 1. Higher levels represent trunk pathways of high-scoring nodes, strongly connected edges, and their associated brain ROIs. We mapped the atlas parcellations of both datasets (i.e., 90 ROIs and 118 ROIs) to Yeo’s seven-network parcellation [48], which includes the visual network (VN), somatomotor network (SMN), dorsal attention network (DAN), ventral attention network (VAN), limbic network (LIM), frontoparietal network (FPN), and default mode network (DMN). Atlas regions not belonging to these networks were classified as ’Others.
@@
Next, we discuss the addiction brain tree illustrated in Figs. 4 (a-1) and (a-2), which features three trunk levels. At the primary trunk level (red pathway), the addiction brain tree predominantly exhibits connectivity among DMN (red), FPN (purple), and partial VAN (sky blue) nodes, similarly, in Figs. 4 (b-1) and (b-2), the schizophrenia brain tree at the main trunk level primarily involves VN (light blue) node, VAN, and partial SMN (green) nodes. At the second hierarchical level (pink pathways), the addiction cohort exhibits predominant connections within DMN and FPN nodes. In contrast, the schizophrenia group at this level demonstrates connectivity primarily among VAN and some DMN nodes. Finally, at the third hierarchical level (light purple pathways), the addiction group’s terminal branches show minimal connections with FPN, VAN, and DMN nodes. In contrast, the schizophrenia cohort displays connections among VN nodes. A key observation is that the addiction brain tree exhibits more distinct communities at different levels, characterized by groups of highly interconnected brain regions. In contrast, the schizophrenia brain tree has a more disordered topological structure, where different networks appear on the same main trunk, leading to a less structured node sequence.
(a) Cannabis
(b) COBRE
Figure 7: Age-related changes in predicted FC values across different groups: (a) cannabis users versus healthy controls (HC) and (b) schizophrenia patients versus HC using the COBRE dataset. Asterisks indicate statistical significance (${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$). Error bars represent the standard error of the mean.
@@
Datasets Age Group Pearson (r) MSE
Cannabis 18-25 (n=150) $0.63\scriptstyle{\pm 0.08}$ $\textbf{7.8}\scriptstyle{\pm 3.1}$
25-40 (n=127) $0.77\scriptstyle{\pm 0.06}$ $15.8\scriptstyle{\pm 5.3}$
40-54 (n=41) $0.72\scriptstyle{\pm 0.05}$ $46.6\scriptstyle{\pm 22.8}$
Overall Prediction $\textbf{0.95}\scriptstyle{\pm 0.01}$ $12.5\scriptstyle{\pm 4.8}$
COBRE 18-25 (n=35) $0.92\scriptstyle{\pm 0.02}$ $\textbf{5.1}\scriptstyle{\pm 1.9}$
25-40 (n=56) $0.77\scriptstyle{\pm 0.16}$ $26.5\scriptstyle{\pm 16.3}$
40-66 (n=61) $0.97\scriptstyle{\pm 0.01}$ $19.6\scriptstyle{\pm 22.0}$
Overall Prediction $\textbf{0.97}\scriptstyle{\pm 0.07}$ $15.5\scriptstyle{\pm 6.0}$
9 Chronological Brain Age Prediction
In this section, we employed NeuroTree to predict mental health disorders and analyze their effects on brain FC changes across different age groups. Participants were stratified into three age-based cohorts, and brain age was predicted in the testing set. First, NeuroTree predicted FC values across age groups. The results indicate that as age increases, the addiction and mental disorders cohorts exhibit substantial age-related changes in FC, indicating that cannabis use and schizophrenia have similar impacts on the brain FC, as illustrated in Fig. 7. Second, we present Table 3, validating NeuroTree predictions of brain age compared to chronological age.
Overall, in NeuroTree age prediction, the cannabis cohort exhibited a lower MSE than the COBRE cohort. Additionally, we calculated the Pearson correlation coefficient (r) between predicted and actual ages, achieving values exceeding 0.95. Notably, for psychiatric disorder predictions, the MSE for the 18–25 age group was substantially lower, with values of 7.8 for cannabis and 5.1 for COBRE. This finding indicates that younger brains exhibit more stable development patterns, higher neuroplasticity, and more consistent characteristics with smaller individual variations. Furthermore, these brains are less impacted by long-term environmental factors. Specifically, the correlation coefficient (r) reached 0.92 for COBRE predictions related to psychiatric disorders in the 18–25 age group. This high accuracy in younger subjects is likely due to their more stable functional connectivity patterns, which are less affected by neurodegeneration. Additionally, our model achieved an MSE of 19.6 and a correlation coefficient of 0.97 when predicting psychiatric disorders among older adults (40–66 years). However, prediction accuracy for this age group requires consideration of additional variables due to the increasing complexity of psychiatric manifestations with advancing age.
@@ 10 Discussion
In section 8, our analysis of the tree-structured topology revealed distinct patterns of brain subnetwork alterations associated with addiction and schizophrenia. These changes were primarily concentrated in networks governing cognitive functions, including the DMN and the decision-making FPN. This network-specific disruption pattern aligns with previous findings [49, 12, 5]. Additionally, cannabis use increases functional connectivity between the DMN and VAN, which substantially correlates with subjective feelings of being ’high’ [50]. Our findings also support previous research indicating that patients with schizophrenia exhibit considerably reduced FC in the left VAN and VN, with these connectivity abnormalities potentially contributing to perceptual and attentional dysfunction [51, 52].
Moreover, we identified age-related FC subnetwork changes associated with mental disorders, as shown in Fig. 6. In addition, alterations primarily originate in the limbic system (LIM) and progressively extend to networks responsible for executive function (FPN), attention control (DAN, VAN), internal cognition (DMN), and motor-related functions (SMN) with increasing age. In contrast, schizophrenia initially affects the frontoparietal network and limbic system (FPN & LIM) before expanding to broader higher-order and sensorimotor networks (VN, DMN, SMN, VAN, DAN). These findings align with Fig. 7, illustrating dynamic FC changes that evolve with age.

### v3

@@ 6 Experiments
In this section, we first evaluate the proposed NeuroTree performance in brain disease classification (section 6.2) and its application to predicting chronological age (section 8). In section 6.4, we analyze the convergence of the spectral norm for $k$-hop connectivity. Finally, we examine how hierarchical brain trees interpret FC patterns between brain regions and their corresponding seven subnetworks in section 7. Detailed experimental settings for NeuroTree are provided in Appendix H.
@@ 6.1 Datasets
We validated two publicly available fMRI datasets —one focusing on cannabis use disorder and the other on schizophrenia. ① Cannabis (Kulkarni et al., 2023): The cannabis dataset comprises fMRI data from two distinct sources. The data were preprocessed from 3-Tesla fMRI acquisitions, and the mean time series for each subject was computed across 90 ROIs using the Stanford atlas parcellation (Shirer et al., 2012). ② COBRE (Calhoun et al., 2012): The Center for Biomedical Research Excellence (COBRE) dataset includes resting-state fMRI data collected from healthy controls and individuals diagnosed with schizophrenia. All MRI data were parcellated into 118 ROI regions using the Harvard-Oxford atlas. We summarized the statistics and demographics of both datasets in Appendix Table 5.
6.2 Qualitative Results
This section is dedicated to evaluating the performance of NeuroTree against state-of-the-art (SOTA) GCN architectures including 1) $Pearson$ GCN 2) $k$-NN GCN 3) GAT (Veličković et al., 2017) 4) BrainGNN (Li et al., 2021) 5) BrainUSL (Zhang et al., 2023) and BrainGSL (Wen et al., 2023) for graph classification tasks. Additionally, we include the tree path-based model PathNN (Michel et al., 2023). The comparative analysis was conducted using 5-fold cross-validation on two distinct datasets, benchmarking NeuroTree against both baseline models and SOTA approaches. The comprehensive experimental results are presented in Table 1.
@@ 6.2 Qualitative Results
Figure 3: Convergence analysis of $\Phi_{k}(t)$ over $k$-hop The spectral norm of $\Phi_{k}(t)$ reveals differential convergence rates across varying $k$-orders among distinct mental disorders, notably demonstrating that cannabis exhibits a steeper convergence gradient compared to COBRE as $\lambda$ increases.
Compare with Brain GCN Models.
@@
We quantitatively assessed brain network construction using various metrics, employing Pearson correlation and KNN as connectivity measures. As shown in Table 1, the GAT model effectively integrates attention mechanisms into fMRI node features, achieving mean AUC scores of 0.72 and 0.67. In contrast, the conventional GCN is limited to learning solely from topological structures. Furthermore, NeuroTree enhances ACC performance in fMRI-based models, improving BrainGNN (ACC $\uparrow$ 23.72%), BrainUSL (ACC $\uparrow$ 12.30%), and BrainGSL (ACC $\uparrow$ 12.30%) on the cannabis dataset. On the COBRE dataset, BrainGNN, BrainUSL, and BrainGSL further improve ACC by 30.00%, 20.37%, and 27.45%, respectively.
Compare with Path-Based Models.
@@
As shown in Table 1, NeuroTree outperforms existing high-connectivity view field-aware graph GCN models, such as the high-order concatenation layer-based MixHop, achieving over 6% average performance improvement in AUC across all metrics on the cannabis dataset. Furthermore, compared to MixHop on the COBRE dataset, NeuroTree not only achieved a 2% increase in AUC, but also demonstrated a mean performance improvement of 6.3% to 6.5% in ACC, Precision, and Recall. Notably, although GPC-GCN can aggregate multi-level connectivity information by considering the concept of paths, it does not significantly improve classification performance on critical connectivity paths. Similarly, the PathNN framework aggregates path information starting from nodes to generate node representations but shows only limited improvements. In contrast, our NeuroTree achieves a 14% improvement in AUC on the cannabis dataset, while PathNN underperforms compared to MixHop on the COBRE dataset. The experimental results indicate that NeuroTree is superior in capturing temporal FC patterns and predicting outcomes in addiction and psychiatric disorders.
@@
Figure 4: The visualization of the brain tree illustrates psychiatric disorders structured into three hierarchical trunk levels. Panels (a-1) and (b-1) mark the most significant nodes along the tree path. The $l$-1 pathways represent regional connectivity, corresponding to the level three brain maps on the right. Panels (a-2) and (b-2) depict the number of connections using color gradients across the hierarchical levels.
6.3 Ablation Study
The Impact of Objective Function and Demographics.
@@ 6.3 Ablation Study
In our experiments, we evaluate the efficacy of CMFC loss integration within our proposed NeuroTree. The quantitative analysis presented in Table 1 indicates modest improvements in the cannabis dataset; however, its implementation in the COBRE dataset achieved an AUC of 0.71, with consistent performance gains exceeding 3% across metrics, particularly in the characterization of mental disorder features. This enhancement stems from combining the loss function with CMFC loss optimization framework that simultaneously suppresses connectivity between dissimilar regions while amplifying patterns among similar neurological areas. Furthermore, our ablation experiments demonstrate that age serves as a critical modulator in AGE-GCN by adapting parameter $\theta$, substantially influencing the classification of changes in brain connectivity.
@@ 6.4 Convergence Analysis
The intuition for analyzing the rate of spectral norm decay with increasing $k$-hop allows us to examine how rapidly information from distant nodes attenuates across the dynamic brain network structure. More detalil explanation can be found in Appendix I. To further examine the convergence speed of $k$-hop FC in various brain disorders, we analyze the spectral norm under the $k=1$ setting for the convergence of $\Phi_{k}(t)$ in Fig. 3 (left). Although the COBRE dataset initially exhibits slower convergence, particularly with the boundary defined around order 15, the overall convergence in the cannabis dataset is faster than in COBRE across different $k$-hop values. Furthermore, when analyzing the convergence of $\Phi_{k}(t)$ under varying $\lambda$ values, as shown in Fig. 3 (middle) and (right), the $\lambda$ values for cannabis are more concentrated and stable. In contrast, COBRE requires larger $\lambda$ values to achieve convergence. The empirical findings demonstrate that examining the spectral norm of $\Phi_{k}(t)$ across various $k$ values effectively reveals the relationship between high-order patterns and decay characteristics of FC in different mental disorders.
@@ 7 Brain Tree Analysis
This section investigates brain disorders by analyzing hierarchical brain networks using decoded fMRI-derived connectomes. NeuroTree deconstructs the fMRI brain network into a tree-structured component graph consisting of a main trunk and radial branches. We identify key regions associated with brain disorders within each trunk by analyzing optimal weighted high-order tree paths. Our analysis highlights the most relevant functional network regions determining critical pathways within the brain connectome tree.
Exploring Hierarchical Regional Patterns in Brain Disorders.
@@
To better understand how embedded brain features vary across different hierarchical levels of brain subnetworks, NeuroTree predicts each ROI score and performs reranking, as illustrated in Fig. 1. Higher levels represent trunk pathways of high-scoring nodes, strongly connected edges, and their associated brain ROIs. We mapped the atlas parcellations of both datasets (i.e., 90 ROIs and 118 ROIs) to Yeo’s seven-network parcellation (Yeo et al., 2011), which includes the Visual Network (VN), Somatomotor Network (SMN), Dorsal Attention Network (DAN), Ventral Attention Network (VAN), Frontoparietal Network (FPN), and Default Mode Network (DMN). Additionally, we incorporated the Subcortical System (SUB) in our analysis to examine the subcortical structures. Atlas regions not belonging to these networks were classified as ’Others.’ More details can be found at Appendix L.2.
@@
Next, we discuss the addiction brain tree illustrated in Fig. 4 (a-1) and (a-2), which features three trunk levels. At the primary trunk level (red pathway), the addiction brain tree predominantly exhibits connectivity among DMN (red) and VN (purple) nodes; similarly, in Fig. 4 (b-1) and (b-2), the schizophrenia brain tree at the main trunk level primarily involves SUB (pink) nodes and partial VN nodes. At the second hierarchical level (pink pathways), the addiction cohort exhibits predominant connections within DAN, DMN, and VN nodes. In contrast, the schizophrenia group at this level demonstrates connectivity primarily among VN nodes and some FPN (light blue) nodes. At the third hierarchical level (yellow pathways), the terminal branches in the addiction group show minimal connections with the SMN nodes (green) and DAN nodes, whereas the schizophrenia cohort displays several connections among DMN nodes.
8 Brain Age Estimation
Comparing estimated fMRI brain age against actual chronological age in individuals with mental disorders can provide valuable insights into the severity and progression of these conditions (Stankevičiūtė et al., 2020). The physiological actual age of the human brain is unknown and cannot be directly measured; however, we can predict brain age through regression tasks on fMRI data using NeuroTree learned fMRI graph embeddings. We hypothesize that there exist certain gaps between the actual chronological age of healthy individuals and the brain age of those with mental disorders. For example, when the gap between model-predicted brain age and actual chronological age is positive, it indicates accelerated brain aging.
To validate model prediction in brain age, we trained NeuroTree on a healthy control cohort and predicted non-healthy subjects to better measure brain age gaps. In Table 2, we demonstrate NeuroTree’s predictive performance across different age groups in two datasets. Notably, the younger age group (18-25 years) exhibited markedly lower Mean Squared Error (MSE) values of 9.51 in the Cannabis dataset and 4.48 in the COBRE dataset compared to other age groups, indicating greater accuracy in brain age prediction for younger individuals. Compared to younger age groups, NeuroTree’s prediction MSE across both mental disorder datasets provides valuable insight, showing that brain age estimation for older individuals may be affected by greater individual variability in brain changes associated with mental disorders after middle age, while younger groups likely exhibit more consistent patterns of FC during neurodevelopment.
Figure 5: The scatterplot shows the gaps between fMRI-predicted brain age and chronological age for healthy control and mental disorder groups.
@@ 8 Brain Age Estimation
Datasets Age Group MSE Pearson (r)
Cannabis 18-25 (n=150) $9.51\scriptstyle{\pm 1.67}$ $0.27\scriptstyle{\pm 0.02}$
25-40 (n=127) $27.32\scriptstyle{\pm 3.15}$ $0.18\scriptstyle{\pm 0.01}$
40-54 (n=41) $25.12\scriptstyle{\pm 1.40}$ $0.18\scriptstyle{\pm 0.00}$
COBRE 18-25 (n=35) $4.48\scriptstyle{\pm 0.42}$
25-40 (n=56) $37.64\scriptstyle{\pm 2.02}$ $0.10\scriptstyle{\pm 0.01}$
40-66 (n=61) $32.64\scriptstyle{\pm 3.43}$ $0.49\scriptstyle{\pm 0.13}$
@@ 9 Discussion
In section 7, our analysis of the tree-structured topology revealed distinct patterns of brain subnetwork alterations associated with addiction and schizophrenia. These changes were primarily concentrated in networks governing cognitive functions, including the DMN, the visual task VN and the decision-making FPN. This network-specific disruption pattern aligns with previous findings (Kleinhans et al., 2020; Kulkarni et al., 2023; Ding et al., 2024b).
Prior studies have established disrupted connectivity patterns in schizophrenia, particularly within the DMN (Ishida et al., 2024; Yang et al., 2022). We identified intermittent abnormal coupling between subcortical regions, especially the connections with the prefrontal cortex and cerebellum (Ding et al., 2024a; Hancock et al., 2023). Furthermore, areas of the visual cortex were correlated with impaired motor coordination and sensorimotor integration (Verma et al., 2023; Zhao et al., 2022), consistent with the high FC regions illustrated in Fig. 4. Additionally, our analysis revealed distinct age-related differences in brain FC between addiction and schizophrenia (Table 2). We further verify our finding in Fig. 5, which illustrates the relationship between predicted brain age and chronological age. The predicted brain age for both mental disorders is lower than the chronological age, particularly in the higher age group. Both conditions appear to affect age-related brain characteristics, leading to an abnormal relationship between brain age and chronological age. Notably, addiction demonstrates a more significant impact on brain development and aging than schizophrenia, which is consistent with our observations in section 8.